OpenStudy (anonymous):

Can someone please show me how to integrate (from 0 to infinity) (1-e^-t)(e^(-st)). Wolfram Alpha shows ln(1+1/s) but it does not show the steps to prove it.

OpenStudy (perl):

do you need help

OpenStudy (perl):

can you post the link to wolfram so i can look at the integral

OpenStudy (anonymous):

that is the problem I need help with! I know the answer should come out to be ln(1+1/s) but I'm unsure how to get it.

ganeshie8 (ganeshie8):

integration by parts should do right ? by the way, this is the integral you need to evaluate for finding laplace transform of \(\large \rm \frac{1-e^{-t}}{t}\) http://www.wolframalpha.com/input/?i=laplace+transform+%281-e%5E-t%29%2Ft also notice \(\large \rm s\gt 0\) for the integral to converge

OpenStudy (anonymous):

Yea I know It's the Laplace transform but I have to show all of the steps and I don't know if I'm screwing it up or what but I can't get integration by parts to work

ganeshie8 (ganeshie8):

@hartnn

hartnn (hartnn):

so LT can't be used ?

ganeshie8 (ganeshie8):

i think it should be okay to use but there is no LT for 1/t and integration by parts doesnt seem to work

hartnn (hartnn):

tried with this ?

hartnn (hartnn):

i wonder whether Differentiation under integral sign method helps to get the answer without using LT

ganeshie8 (ganeshie8):

nope, i am still messing with by parts but its hopeless really

ganeshie8 (ganeshie8):

\[\begin{align} I(s) &= \int\limits_0^{\infty} \dfrac{1-e^{-t}}{t}e^{-st} \; dt\\~\\ I'(s)&= \int\limits_0^{\infty} \dfrac{\partial}{\partial s}\dfrac{1-e^{-t}}{t}e^{-st} \; dt \\~\\ &= \int\limits_0^{\infty} -s(1-e^{-t})e^{-st} \; dt \\~\\ &= s\int\limits_0^{\infty} -e^{-st} +e^{-t(s+1)} \; dt \\~\\ & = s\left[\dfrac{e^{-st}}{s} - \dfrac{e^{-t(s+1)}}{s+1}\right]_{t=0}^{t=\infty} \\~\\ &=-\dfrac{1}{s+1} \end{align}\] integrate both sides and thank @hartnn :)

ganeshie8 (ganeshie8):

small mistake :\[\begin{align} I(s) &= \int\limits_0^{\infty} \dfrac{1-e^{-t}}{t}e^{-st} \; dt\\~\\ I'(s)&= \int\limits_0^{\infty} \dfrac{\partial}{\partial s}\dfrac{1-e^{-t}}{t}e^{-st} \; dt \\~\\ &= \int\limits_0^{\infty} -(1-e^{-t})e^{-st} \; dt \\~\\ &= \int\limits_0^{\infty} -e^{-st} +e^{-t(s+1)} \; dt \\~\\ & = \left[\dfrac{e^{-st}}{s} - \dfrac{e^{-t(s+1)}}{s+1}\right]_{t=0}^{t=\infty} \\~\\ &=-\dfrac{1}{s}+\dfrac{1}{s+1} \end{align}\]

OpenStudy (anonymous):

is the final answer that above. Should there be an ln in there somewhere?

ganeshie8 (ganeshie8):

you have \[\large \rm I'(s) = -\dfrac{1}{s} + \dfrac{1}{s+1}\] integrate both sides for\(\large \rm I(s)\)

OpenStudy (anonymous):

ahhhhhh ok thanks!!!!!

ganeshie8 (ganeshie8):

np:) go thru that MSE link if u have time... it has half dozen different methods for evaluating this integral